OpenStudy (anonymous):

When m is divided by 14, the remainder is 7. When is the remainder is 7. What is the remainder when 9m is divided by 14?

hartnn (hartnn):

is your question complete ??

OpenStudy (anonymous):

yea

hartnn (hartnn):

in this case u need to use the formula \(\text{Dividend=Divisor}\times \text{quotient+remainder}\)

OpenStudy (anonymous):

@hartnn i think when is the remainder should not b there. @Soni-kay from the first sentence m = 14*x +7, 9m= 14*9x +7*9 = 14*9x + 63 so find remainder when 63 is divided by 14, that will be your answer

OpenStudy (anonymous):

7...if u divide any no. n get a remainder...the remainder remain same if the no. is multiplied n times..and divisor remain same...

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

@pkjha3105 , if i divide 3 by 2, reminder is 1, now i multiply 3*2, now the reminder is 0, by your arguement this should be 1 right